I have another difficult problem
x^2/3-3x ^1/3-4=0

Answer by Alan3354(69443)

Sub u for x^(1/3)
----------
u^2 - 3u - 4 = 0
(u-4)*(u+1) = 0
u = -1, 4
-------------
x^(1/3) = -1
x = -1
----------
x^(1/3) = 4
x = 64
Not that hard, just a bit tricky.